OpenStudy (anonymous):

Find a polar equation for the conic described: The ellipse with a focus at the pole and major axis endpoints of (5,0) and (3, pi)

OpenStudy (anonymous):

Cannot draw. Tips when center of mass is no longer above support. E.g. square center would be just beyond the horizontal position of its corner.

OpenStudy (anonymous):

Huh?....Not a clue what you mean.

OpenStudy (anonymous):

Supposed to fit one of the polar forms, though.

OpenStudy (zale101):

oh ok, then nvm sorry

OpenStudy (anonymous):

Like \[\frac{ ep }{ 1 \pm esin \theta } or \frac{ ep }{ 1 \pm ecos \theta }\]

OpenStudy (zale101):

i sent u a link, i think it might help u

OpenStudy (anonymous):

Glancing it over right now.

OpenStudy (anonymous):

Yeah, not enough information to help me out much x_x

OpenStudy (zale101):

@ganeshie8

ganeshie8 (ganeshie8):

@ikram002p

ganeshie8 (ganeshie8):

lets try, (5,0) and (3, pi) major axis on 0, pi => its a horizontal ellipse

OpenStudy (anonymous):

Right, I can gather that much, but I dont know how to use that to get e or p or even determine whether it is sin or cos in the equation. Im guessing cos since the angle pi gave a value of 3, but not sure. \[\frac{ ep }{ 1 \pm \cos \theta }?\]

OpenStudy (anonymous):

Oops \[r = \frac{ ep }{ 1 \pm ecos \theta }\]

ganeshie8 (ganeshie8):

directrix is vertical => its a cos \(\large r = \frac{ep}{1+ e \cos \theta }\)

ganeshie8 (ganeshie8):

we need to find e and p

OpenStudy (anonymous):

How can you tell its vertical?

ganeshie8 (ganeshie8):